OpenStudy (anonymous):

express 3^x, using the base e sorry lol i havent done this math in awhile and my little sister needs help

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (anonymous):

You need to understand logarithms.

OpenStudy (anonymous):

\[e^{\ln 3} = 3\]

OpenStudy (anonymous):

So \[ (e^{\ln 3})^x = 3^x \]

OpenStudy (anonymous):

Or basically \[ 3^x = e^{x\ln 3} \]